OpenStudy (anonymous):

what is the value of 5k to the power of 2/ - 6 when k = 9? A. 84 B. 375 C. 399 D. 2019

OpenStudy (jango_in_dtown):

kk^(2-6)??? .@johnnyy999

OpenStudy (anonymous):

5k^2-6 k=9

OpenStudy (michele_laino):

you have to solve this quadratic equation: \(5k^2-6k-9=0\)

OpenStudy (jango_in_dtown):

itss 399 since 5(9^2)-6=5*81-6=405-6

OpenStudy (jango_in_dtown):

=399

OpenStudy (anonymous):

yeah! @Michele_Laino

OpenStudy (michele_laino):

oops.. It is: \(5 \cdot 9^2- 6=...\)

OpenStudy (anonymous):

yeah dude...

OpenStudy (anonymous):

anyone know the answer?

OpenStudy (michele_laino):

I'm sorry, I misunderstood: it is not a quadratic equation

OpenStudy (jango_in_dtown):

399 i gave it

OpenStudy (anonymous):

soooo... C are u shure??

OpenStudy (jango_in_dtown):

yeah /..

OpenStudy (jango_in_dtown):

you calculate and check

OpenStudy (anonymous):

ok thx...

OpenStudy (michele_laino):

that's right: it is \(399\): \(5 \cdot 81 -6= 405-6=399\)